# MOTI: A Drug-Target Interaction Graph For Inductive Link Prediction

### v1

@@ 4.2 Inductive link prediction benefits from graph structure and CP features
| GraphSAGECP | 0.43400.009 | 0.1721 0.009 | 0.78160.016 | |
| Source | MLP | 0.1914 0.018 | 0.0217 0.008 | 0.1700 0.052 |
| Bilinear | 0.1681 0.000 | 0.0109 0.000 | 0.0988 0.003 | |
| GraphSAGECP | 0.3903 0.011 | 0.10250.004 | 0.54400.005 | |
| Target | MLP | 0.2118 0.007 | 0.0214 0.005 | 0.1736 0.030 |
| Bilinear | 0.1702 0.001 | 0.0159 0.001 | 0.1280 0.010 |
We evaluated the models on inductive link prediction tasks using the cold data splits, which require informative node features as the left out nodes are completely isolated during evaluation [5]. GraphSAGEembs and the shortest path heuristic are not applicable here, as GraphSAGEembs cannot represent new, unseen entities and shortest path cannot handle isolated nodes unreachable by any path.
For both cold-source and cold-target split, GraphSAGECP greatly outperformed the feature-based models. For all three models, the only accessible information about the left out nodes is their feature vectors, as they are isolated and without edge connections. Therefore, the improved predictions by GraphSAGECP can be attributed to it having learned improved source and target node embeddings by leveraging both the node features and the relational information present in the graph structure during training.
Inductive link prediction is useful in real world applications as the model is learning about and making predictions on entities that we previously knew very little about. The cold-source split is especially relevant for drug discovery, as it simulates the scenario where all genes are known and the DTI model is tasked with identifying new compounds associated with certain disease-related genes. Furthermore, empirical features, such as morphological profiles, prove to be valuable node representations. If we collect experimental data for a newly discovered compound or unexplored gene, GraphSAGECP and similar models would be able to leverage known relationships between other compounds and genes to make better inductive link predictions for the new entity. This underscores the value of integrating assay-based data into graph-based models to enhance their predictive capabilities.
@@ 4.3 Ablation studies with graph structure
We investigated the contribution of each edge type (source-target, source-source, and target-target) by comparing the performances of the GraphSAGE models using the four graph structures (Figure 2). We see that GraphSAGECP outperforms GraphSAGEembs for every graph structure across every metric. As we add more edge types to the graph, the performance difference between GraphSAGECP and GraphSAGEembs diminishes. This suggests that the node features have a more significant impact on the model’s performance when the graph structure is sparse. As the graph becomes denser with the addition of more edge types, the structural information begins to dominate, reducing the relative contribution of the node features. These findings highlight the complementary nature of node features and graph structure in the link prediction task.
@@ 4.4 Zero-shot prediction potential
To assess the model’s zero-shot prediction potential, we computed the average precision scores for each source and target node in the test set of a cold-source split run. In this scenario, all source nodes in the test set are completely unseen during training, and most but not all target nodes are seen (a target node may not be seen if it is only connected to left out sources). Thus, the model’s prediction of an edge with isolated endpoints, i.e. both source and target are unseen during training, is a zero-shot link prediction. As shown in Figure 3, the presence of unseen target nodes with high average precision scores (rightmost blue bar in target plot) indicates that GraphSAGECP can make accurate predictions for novel source-target pairs. This suggests the model is able to generalize knowledge learned from the training data to make informed predictions in the zero-shot setting, motivating further research in this direction.
@@ 5 Discussion
Predicting the complex relationships between chemical compounds and genes is ambitious. Developing models to accurately predict these high-order interactions, particularly in inductive or zero-shot settings, is even more challenging. The MOTI dataset addresses these challenges by integrating morphological features of cells and known compound-gene relationships. By providing a large-scale, feature-rich, and extensively annotated dataset, MOTI enables the development and benchmarking of graph-based models for DTI prediction in transductive, inductive, and zero-shot scenarios. One of the key strengths of MOTI lies in its rigor. The morphological profiles were extracted from the JUMP CP dataset using a standardized pipeline, ensuring consistency and reproducibility. The careful curation of annotations prioritized data quality and reliability. The constructed dataset also offers stringent protection against data leakage of any kind, while providing thorough and challenging forms of data splitting, loading, minibatch training (requiring subgraphs), sampling, and evaluation procedures.
We acknowledge certain limitations in MOTI. As discussed in Section 4.3, the effectiveness of CP features may be reduced when the graph is densely connected. Additionally, some perturbations may not induce significant morphological changes, potentially leading to uninformative node representations and false positive predictions. To address these issues, we propose isolating samples with distinguishable morphologies and reanalyzing them to assess the impact on performance.
We only utilized ORF profiles for target nodes but future work may incorporate both the ORF and CRISPR gene features as multimodal inputs in the graph. Additionally, methods may be extended to predict heterogeneous interactions rather than just binary classifications. Finally, developing end-to-end architectures to learn node embeddings directly from images could better exploit the morphological information. MOTI represents a valuable resource for the machine learning community, particularly for those interested in graph-based methods and their applications in drug discovery. By fostering interdisciplinary collaborations across graph ML, biological imaging, and drug discovery, MOTI has the potential to accelerate progress for the challenging and complex task of DTI prediction.
@@ Appendix A GraphSAGECP forward pass algorithm
| Convolution weight matrices and |
@@ Appendix B Experimental Set Up
We performed a random hyperparameter search [31] for each model and data split for the number of hidden channels (), learning rate (), and weight decay (). We sampled negative edges at a ratio of 1:1 at training and validation time, and 1:10 at test time. We trained each model for 1000 epochs, computed the Binary Cross Entropy loss between the prediction scores for the supervision edges and their ground truth labels, and used an Adam Optimizer to make weight updates. At validation and test time, we computed the F1 scores of the predicted edges using the best threshold found during validation time. We also computed two rank-based metrics, Hits@500 and Precision@500, to better capture how the model distinguishes between positive and negative samples. Hits@ quantifies the fraction of positive test edges (total = ) that rank () within the top negative test edge scores 222https://ogb.stanford.edu/docs/linkprop/. Precision@ quantifies the fraction of the top predicted scores that are assigned to true positive edges. In both cases, was chosen to isolate the around of the test edge scores.
@@
We ran each scenario 5 times, and we reported the average performance for each model as well as the standard deviation scores in Tables 3 and 4. We set a random seed for all runs and established the same data split for each of our splitting methods, such that all variance in performance must come from batch sampling by the data loader and GPU non-determinism. All experiments ran in a NixOS server with a AMD Ryzen Threadripper PRO 7995WX processor and an NVIDIA RTX 6000 GPU card.

### v2

@@ 4.2 Inductive link prediction benefits from graph structure and CP features
GINCP | 0.28270.018 | 0.10780.008 | 0.58720.028 | |
GraphSAGECP | 0.22830.014 | 0.04390.013 | 0.31200.079 | |
Source | GATv2CP | 0.24330.009 | 0.03300.002 | 0.24760.004 |
MLP | 0.18760.006 | 0.02690.009 | 0.21680.063 | |
Bilinear | 0.15930.004 | 0.01450.003 | 0.12560.029 | |
GINCP | 0.39160.016 | 0.27440.020 | 0.96240.034 | |
GraphSAGECP | 0.34940.012 | 0.19880.013 | 0.70560.031 | |
Target | GATv2CP | 0.28050.007 | 0.19660.019 | 0.95400.043 |
MLP | 0.34080.014 | 0.18370.010 | 0.65520.025 | |
Bilinear | 0.13960.001 | 0.01220.001 | 0.10240.004 |
We evaluated the models on inductive link prediction tasks using the cold data splits, which require informative node features as the left out nodes are completely isolated during evaluation [5]. In this scenario, GraphSAGEembs is not applicable as it learns representations for nodes using projection layers and is thus limited to making predictions for nodes that it has already seen during training (i.e. transductive link predictions). The shortest path heuristic is also not applicable, as it is unable to handle isolated nodes unreachable by any path.
During inductive link prediction, all of the models only have access to the feature vectors of left-out nodes, as they are isolated and without edge connections. For both cold-source and cold-target split, GINCP greatly outperformed all other models. The improved predictions by GINCP can be attributed to it having learned improved source and target node embeddings by leveraging both the node features and the relational information present in the graph structure during training. Except for the F1 score of the GATv2 model in the cold-target split, all of the GNN models outperform the feature-based models. This supports our hypothesis that knowing the relationships between sources and targets improves DTI prediction.
Inductive link prediction is useful in real world applications as the model is learning about and making predictions on entities that we previously knew very little about. The cold-source split is especially relevant for drug discovery, as it simulates the scenario where all genes are known and the DTI model is tasked with identifying new compounds associated with certain disease-related genes. Furthermore, empirical features, such as morphological profiles, prove to be valuable node representations. If we collect experimental data for a newly discovered compound or unexplored gene, GINCP and similar models would be able to leverage known relationships between other compounds and genes to make better inductive link predictions for the new entity. This underscores the value of integrating assay-based data into graph-based models to enhance their predictive capabilities.
@@ 4.3 Ablation studies with graph structure
We investigated the contribution of each edge type (source-target, source-source, and target-target) by comparing the performances of the GraphSAGE models using the four graph structures (Figure 2). While this analysis can apply to any of the GNN models, we chose GraphSAGE such that we could compare the feature-based and embeddings-based cases. We see that GraphSAGECP outperforms GraphSAGEembs for almost every graph structure and metric. More analyses should be done to investigate how the difference in performance fluctuates with the addition of edge types in order to understand how the usefulness of the node features fluctuates with graph sparsity. From these initial results, it appears that as the graph becomes denser with the addition of more edge types, the structural information begins to dominate, reducing the relative contribution of the node features. These findings highlight the complementary nature of node features and graph structure in the link prediction task.
@@ 4.4 Zero-shot prediction potential
To assess the model’s zero-shot prediction potential, where neither the source nor the target nodes are seen in training or validation, we computed the average precision (AP) scores for each source and target node to evaluate how well the model was able to retrieve its true links. In this cold-source split scenario, all source nodes in the test set are completely unseen during training, and most but not all target nodes are seen (a target node may not be seen if it is only connected to left out sources). Thus, the model’s prediction of an edge with isolated endpoints, i.e. both source and target are unseen during training, is a zero-shot link prediction. As shown in Figure 3, the presence of unseen target nodes with high AP scores (see the rightmost blue bar in the Target panel) shows that for some node pairs, the model is able to predict many true links even though both the source and target nodes are isolated at test time. This suggests the model, trained on CP features and known drug-gene relationships, can predict links between completely novel drugs and genes. Further work should be done to explore the efficacy of DTI predictions in the zero-shot setting.
@@ 5 Discussion
Predicting the complex relationships between chemical compounds and genes is ambitious. Developing models to accurately predict these high-order interactions, particularly in inductive or zero-shot settings, is even more challenging. The MOTI dataset addresses these challenges by integrating morphological features of cells and known compound-gene relationships. By providing a large-scale, feature-rich, and extensively annotated dataset, MOTI enables the development and benchmarking of graph-based models for DTI prediction in transductive, inductive, and zero-shot scenarios. One of the key strengths of MOTI lies in its rigor. The morphological profiles were extracted from the JUMP CP dataset using a standardized pipeline, ensuring consistency and reproducibility. Carefully curating annotations prioritized data quality and reliability. The constructed dataset also offers stringent protection against data leakage of any kind while providing thorough and challenging forms of data splitting, loading, minibatch training (requiring subgraphs), sampling, and evaluation procedures.
We acknowledge certain limitations in MOTI. As discussed in Section 4.3, the effectiveness of CP features may be reduced when the graph is densely connected. Additionally, some perturbations may not induce significant morphological changes, potentially leading to uninformative node representations and false positive predictions. To address these issues, we propose isolating samples with distinguishable morphologies and reanalyzing them to assess the impact on performance. Finally, we recognize that the availability of image-based profiles remains a limitation for expanding our DTI graph; we propose that MOTI could be extended by training a generative model that translates the compound structure to in-silico Cell Painting readouts, as similarly explored in Zapata et al. [35]. Also, if the network is extended to include multiple modalities, then it could also be adapted to make predictions for nodes with missing modalities [36].
Future work may incorporate both the ORF and CRISPR gene features as a form of multimodal inputs in the graph. MOTI could also be expanded to include alternative representations for compound and gene nodes (e.g. protein structures in addition to image-based profiles) to capture known structural and sequencing relationships. The complementary information between profiles may then lead to higher quality DTI predictions. Additionally, methods may be extended to predict heterogeneous interactions rather than just binary classifications. Finally, developing end-to-end architectures to learn node embeddings directly from images could better exploit the morphological information. MOTI represents a valuable resource for the machine learning community, particularly for those interested in graph-based methods and their applications in drug discovery. By fostering interdisciplinary collaborations across graph ML, biological imaging, and drug discovery, MOTI has the potential to accelerate progress for the challenging and complex task of DTI prediction.
@@ Appendix A GraphSAGECP forward pass algorithm
Message Passing Network and |
@@ Appendix B Experimental details
We performed a random hyperparameter search [37] for each model and data split for the number of hidden channels (), learning rate (), and weight decay (). We sampled negative edges at a ratio of 1:1 at training and validation time, and 1:10 at test time. We trained each model for 1000 epochs, computed the Binary Cross Entropy loss between the prediction scores for the supervision edges and their ground truth labels, and used an Adam Optimizer to make weight updates. At validation and test time, we computed the F1 scores of the predicted edges using the best threshold found during validation time. We also computed two rank-based metrics, Hits@500 and Precision@500, to better capture how the model distinguishes between positive and negative samples. Hits@ quantifies the fraction of positive test edges (total = ) that rank () within the top negative test edge scores 222https://ogb.stanford.edu/docs/linkprop/. Precision@ quantifies the fraction of the top predicted scores that are assigned to true positive edges. In both cases, was chosen to isolate the around of the test edge scores. To define these metrics more formally, let represent the rank of the -ranked negative edge. Then, the following equations apply:
@@
We ran each scenario five times, and we reported the average performance for each model as well as the standard deviation scores in Tables 3 and 4. We set a random seed for all runs and established the same data split for each of our splitting methods, such that all variance in performance must come from batch sampling by the data loader and GPU non-determinism. All experiments ran in a NixOS server with a AMD Ryzen Threadripper PRO 7995WX processor and an NVIDIA RTX 6000 GPU card.
Checklist
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
Did you describe the limitations of your work? [Yes] In Section 5
Did you discuss any potential negative societal impacts of your work? [No] As MOTI relies on publicly available compound-gene relationships and morphological profiles and does not involve any human subject data, we do not believe this dataset may be misused to have a negative societal impact.
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] In the abstract and at https://github.com/carpenter-singh-lab/motive
Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] In Appendix B
- (c)
Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] In Appendix B
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
If your work uses existing assets, did you cite the creators? [Yes] We compiled existing databases and cited the datasets in Table 1
Did you mention the license of the assets? [No] We will discuss the licensing of dependency assets and our asset in the dataset documentation as part of the supplementary materials.
Did you include any new assets either in the supplemental material or as a URL? [Yes] We will include detailed directions for how to access our dataset in the supplementary materials. The abstract also links to the GitHub page, where the asset can be accessed.
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] All of the data we used and curated was publicly available and published online.
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [No] We believe we made clear from the description of our dataset that it does not contain personally identifiable information or offensive content.
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]